# Not Every AI Problem is a Data Problem: We Should Be Intentional About Data Scaling

### v1

@@ Abstract
While Large Language Models require more and more data to train and scale, rather than looking for any data to acquire, we should consider what types of tasks are more likely to benefit from data scaling. We should be intentional in our data acquisition. We argue that the topology of data itself informs which tasks to prioritize in data scaling, and shapes the development of the next generation of compute paradigms for tasks where data scaling is inefficient, or even insufficient.
@@
Scaling laws for LLMs have been driving our thinking on how to build optimal models almost since the inception of transformers (Vaswani, 2017). In 2020, OpenAI researchers demonstrated the relationships between compute budget, model size, and dataset size for an optimal model (Kaplan et al., 2020). Subsequently, researchers from Google DeepMind investigated the compute-optimal growth of model and dataset size under a fixed compute budget, demonstrating that the model size and training data should be scaled equally (Hoffmann et al., 2022). The scaling narrative suggests that this path could continue to yield improvements indefinitely, allowing LLMs to get better and better, and addressing more and more use cases over time.
@@
Consider Machine Translation (MT). Early deterministic AI translation models struggled with contextual nuances of language (American Translators Association, 2018), such as capturing culturally specific elements. Now transformer-based language models excel in this area due to several factors. First, the relatively static nature of language, with its abstractable rules and gradual vocabulary evolution, offered a stable foundation for model training. Second, high quality translation data, often sourced from reputable publishers and professional translations, further enabled effective training. Expansive datasets, including non-parallel corpora, enhanced the naturalness and contextual understanding of these models. Finally, LLM’s vastly superior capacity to understand contexts (e.g., an entire document) led to significant improvements in translation quality.
@@
We should be intentional in data-driven scaling. By focusing on use cases with a strong hypothesis about efficacy of scaling, and by collecting fit-for-purpose data based on the needs of these use cases, we can make model training more efficient and reduce the volume of data that we need. Even with existing data, intentional filtering and selection are crucial to ensure that a larger fraction of training mixture is of high quality.
@@ Acknowledgments
We are extremely grateful to Andrew Dai, Anoop Sinha, Blaise Aguera y Arcas, Isabelle Guyon, Rahul Sukthankar, Ethan Dyer, Rosanne Liu, Kenric Allado-McDowell, and Brendan Conway-Smith for their input and discussions.

### v2

@@ Abstract
While Large Language Models require more and more data to train and scale, rather than looking for any data to acquire, we should consider what types of tasks are more likely to benefit from data scaling. We should be intentional in our data acquisition. We argue that the shape of the data itself, such as its compositional and structural patterns, informs which tasks to prioritize in data scaling, and shapes the development of the next generation of compute paradigms for tasks where data scaling is inefficient, or even insufficient.
@@
Scaling laws for LLMs have been driving our thinking on how to build optimal models almost since the inception of transformers (Vaswani, 2017). In 2020, OpenAI researchers demonstrated the relationships between compute budget, model size, and dataset size for an optimal model (Kaplan et al., 2020). Subsequently, researchers from Google DeepMind Hoffmann et al. (2022) and other frontier labs investigated the compute-optimal growth of model and dataset size under a fixed compute budget, demonstrating that the model size and training data should be scaled equally. The scaling narrative suggests that this path could continue to yield improvements indefinitely, allowing LLMs to get better and better, and addressing more and more use cases over time.
@@
Consider Machine Translation (MT). Early deterministic AI translation models struggled with contextual nuances of language (American Translators Association, 2018), such as capturing culturally specific elements. Now transformer-based language models excel in this area due to several factors. First, the relatively static nature of language, with its abstractable rules and gradual vocabulary evolution, offered a stable foundation for model training. Second, high quality translation data, often sourced from reputable publishers and professional translations, further enabled effective training. Expansive datasets, including non-parallel corpora, enhanced the naturalness and contextual understanding of these models. Finally, LLM’s vastly superior capacity to understand contexts (e.g., an entire document) led to significant improvements in translation quality. As a result, LLM-based techniques have brought such significant improvements to machine translation that they have become the new standard Ciesielski (2024).
Robotics—from autonomous driving to factory automation—also demonstrates how data-driven scaling can outperform purely rule-based approaches. Vast sensor logs capture many of the scenarios and environments where robots operate. The “long tail” of rare events, such as extreme conditions or erratic human behavior, demands richer inputs. Higher-resolution imagery and synthetic data increasingly address these demands. As sensor and camera costs drop and real-world data becomes more abundant, these conditions create a virtuous cycle of continuous improvement and propel robotics toward greater reliability than what was possible with prior methods.
Similarly, drug discovery benefits from a vast trove of past experiments, while fundamental biological processes remain relatively stable over time. However, coverage gaps, such as underexplored compound families, pose significant challenges. Pinpointing these gaps ensures that newly gathered data remains relevant and interpretable for model training. Meanwhile, stable features that emerge across experiments highlight critical research directions, enabling more targeted and intentional data acquisition.
@@
Figure 1: Operationalizing data-shape based assessment of use cases: The table lists the types of questions that will help understand the shape of data for specific use cases. We demonstrate the application of this framework to two use cases, machine translation and understanding misinformation.
@@
We should be intentional in data-driven scaling. By focusing on use cases with a strong hypothesis about efficacy of scaling, and by collecting fit-for-purpose data based on the needs of these use cases, we can make model training more efficient and reduce the volume of data that we need. In turn, we will make model training more efficient and sustainable. Even with existing data, intentional filtering and selection are crucial to ensure that a larger fraction of training mixture is of high quality.
@@ Acknowledgments
We are extremely grateful to Jennifer Prendki, Andrew Dai, Anoop Sinha, Blaise Aguera y Arcas, Isabelle Guyon, Rahul Sukthankar, Ethan Dyer, Rosanne Liu, Kenric Allado-McDowell, and Brendan Conway-Smith for their input and discussions.